Describe how would you solve the following problem using an algebraic equation: The perimeter of the rectangle is 24 ft. If the length is three times as long as the width what is the length of the rectangle?
Answer by CharlesG2(834)

Describe how would you solve the following problem using an algebraic equation: The perimeter of the rectangle is 24 ft. If the length is three times as long as the width what is the length of the rectangle?
Perimeter P rectangle = 2 * (L length + W width) = 24 ft
P = 2(L+W) = 24
L + W = 12 (divided both sides by 2)
L = 3W
3W + W = 12 (used substitution)
4W = 12
W = 3 ft so L = 9 ft